Optimization

Find the dimensions of a rectangle with a perimeter of 76 feet that has the maximum area.

Find the dimensions of a rectange worth a primitir. 876 feet freut han the naikinillnearice.
Answer #1

let of then, a xb be the dimensions rectangle perimeter of rectangle P = 2 (a+b) = 78 => a+b = 39 Now, Area b = 39-a of recta